OpenStudy (anonymous):

how do you derive y=x^(1/2)?

OpenStudy (anonymous):

memorize this one it is very common \[\frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt{x}}\] don't forget it

OpenStudy (anonymous):

Of course, if u do happen to forget it notwithstanding satellite's good advice, you can just do it normally (power rule or whatever they call it now) 1/2 x^(-1/2))

OpenStudy (anonymous):

yes of course. but while you are taking an exam or doing homework you really want to just write the answer. kind of like memorizing 8 times 7, whatever that is